# Physics beyond the Standard Model with the DSA-2000

### v1

@@ 2.4 Axion clouds around neutron stars
Fig. 7 shows the projected reach for DSA-2000 observation of the Crab pulsar (dashed magenta), ATNF catalog (solid green), and the new population model (solid blue), requiring an SNR of $5$. We also include the projected reach for a $10$-hr observation of the best target predicted by our population model to be discovered by the DSA-2000 (dashed red), as well as an optimistic $1000$-hr observation of the same target (dotted red). Existing constraints in this mass range are shown gray. Leading astrophysical constraints on non-DM axion (darker gray) include the radio signal from polar cap axion conversion to photon exiting the NS [106]. Leading terrestrial constraints on axion DM (lighter gray) includes ADMX [80, 81, 82, 83, 84, 85, 86], CAPP [91, 92, 93, 94, 95, 96, 97, 98, 99, 100], and RBF+UF [87, 88, 89, 90]. Our projections indicate that DSA-2000 will reach a sensitivity of order $g_{a\gamma\gamma}\sim 10^{-14}\,\text{GeV}^{-1}$, which will set the most competitive non-DM axion constraints in the mass range $2.9\,\mu\mathrm{eV}\lesssim m_{a}\lesssim 8.3\,\mu\mathrm{eV}$, including the uncovered parameter space in the mass range $5\,\mu\mathrm{eV}\lesssim m_{a}\lesssim 8\,\mu\mathrm{eV}$ in the axion DM landscape. This can be achieved with either a $10$-hr fixed-target search pointing at the Crab pulsar, or data from the planned all-sky survey assuming discovery of the expected new pulsars. Independent of the discovery of new pulsars, data from the planned all-sky survey will allow us to set the most competitive non-DM axion bound at $\sim 10^{-13}\,\text{GeV}^{-1}$ in the mass range $3\,\mu\mathrm{eV}\lesssim m_{a}\lesssim 8\,\mu\mathrm{eV}$, using the ATNF catalog. If DSA-2000 is able to discover a pulsar with parameters comparable to the best predicted target (see Table 3), a $10$-hr dedicated observation would allow us to reach $g_{a\gamma\gamma}\approx 5\times 10^{-15}\,\text{GeV}^{-1}$, and a $1000$-hr dedicated observation would probe the KSVZ axion hypothesis.
@@ 3 Radio signals from dark photons
Many of the observational signatures of axions are shared by another well-motivated DM candidate, the dark photon (sometimes referred to as the “hidden photon”) [132, 133, 134, 135]. Indeed, resonance conversion into radio photons near NSs is also possible for dark photon DM [136], in close analogy with the axion DM case discussed in Sec. 2. However, the limits one could obtain from such radio signals in the DSA-2000 frequency band are not competitive. Here we focus on using the DSA-2000 to probe radio signals generated through dark photon superradiance [2], with sensitivity to dark photon masses that range from $5\times 10^{-14}-10^{-12}\,\text{eV}$.
The dark photon, $\gamma^{\prime}$, is a hypothetical vector particle with mass $m_{\gamma^{\prime}}$ that interacts with the Standard Model via kinetic mixing with ordinary photons. This is a well-motivated extension to the Standard Model that could, for example, arise from a new $U^{\prime}(1)$ gauge group. The dark photon itself is a viable DM candidate, but could also constitute a vector portal to a dark BSM sector which contain the stable DM. It is extensively searched for using laboratory probes, for example, fifth force [137, 138] and light shining through wall experiments [139, 140], astrophysical probes [141, 142], as well as cosmological probes such as cosmic microwave background measurements (CMB)[143, 144, 145]. Its interaction with the SM is described by the Lagrangian
@@
where $A^{\mu}$ denotes the SM photon, $F^{\mu\nu}$ is the electromagnetic field strength, $A^{\prime\mu}$ is the dark photon, $F^{\prime\mu\nu}$ is the dark photon field strength, $\epsilon$ is the kinetic mixing parameter, and $J^{\mu}$ ($J^{\prime\mu}$) is the electromagnetic (dark) current. While a priori the mixing parameter $\epsilon$ could be order one, experimental probes constrain it to be much smaller. Dark photons can oscillate during propagation in vacuum or matter from their sterile state to visible photons. This oscillation can be resonantly enhanced if the dark photon mass, $m_{\gamma^{\prime}}$, coincides with the plasma mass of the visible photon, which scales with the electron number density. This process, analogous to resonant axion-photon conversion in plasma, sets the most stringent constraints on non-DM dark photons, $\epsilon\lesssim 5\times 10^{-7}-5\times 10^{-8}$ on dark photons in the relevant mass range from $10^{-14}-10^{-11}\,\text{eV}$ from measurements of spectral distortions [146, 145] and spectral anisotropies [147, 148, 149] in the CMB. [^9] We disregard the constraints from [150] who argue that black hole superradiance rules out dark massive vector particles in the mass range from $10^{-13}-3\times 10^{-12}\,\text{eV}$ based on the observed stability of the black holes’ inner disk. Non-gravitational interactions of the cloud perturb the disk dynamics as explored in [2] which invalidates an extrapolation of the result of [150] to the finite values of $\epsilon$ we consider.
@@ 3.1 Kinetically mixed dark photon superradiance
The phenomenon of Black Hole (BH) superradiance describes the formation of a bosonic cloud around a spinning black hole [151, 152, 153, 154, 155, 156]. The bosonic cloud is a coherent gravitationally-bound state of ultralight bosons, which forms if an ultralight boson exists whose Compton wavelength is of the order of the BH horizon size [157, 158, 156]. This phenomenon is independent of any initial abundance such that superradiance occurs even if the boson does not comprise any sub-component of DM. For a dark photon, the superradiance mechanism sources a dark photon superradiance cloud around a spinning BH [159], which if it possesses a large enough kinetic mixing with the SM photon, induces a rotating visible electromagnetic field, accompanied by a plasma of charged particles [2]. The number of dark photons in the cloud grows exponentially, following the birth of a BH with initial dimensionless spin $a_{*}$, mass $M$, and angular velocity [^10] The angular velocity of a BH indicates how fast its event horizon is rotating.[160, 2]
@@
Here $G$ denotes Newton’s gravitational constant, and [2]
@@
The conducting plasma within the superradiance cloud is seeded by initial synchrotron radiation from environmental electrons accelerated in the EM field of the superradiance cloud [2]. They induce photon-assisted Schwinger pair production, efficiently populating an EM plasma for kinetic mixing parameters that exceed about [2]
@@
[…]
@@
Unlike normal pulsar signals, the time derivative of the pulsing frequency is positive, indicating a spin-up, [2]
@@
Since the pulsing frequency at leading order is determined by the dark photon mass, identifying several objects with these characteristics clustered around the same frequency would be a smoking gun for detecting a kinetically mixed dark photon superradiance cloud, particularly if the signals can be spatially correlated with prior BH merger events. The characteristic dark photon masses that would lead to this observable superradiance phenomenon are around $m_{\gamma^{\prime}}\sim 10^{-12}\,\text{eV}$ for stellar BHs [2], which corresponds to dark photons whose Compton wavelengths are of the order of the observed BH horizon sizes.
DSA–2000 can probe this dark photon signal through different search strategies. In particular, electromagnetic follow-ups of compact binary mergers, with characteristics that allow for a timely formation ($t_{\text{growth}}\lesssim\mathcal{O}\,(\text{years})$) of the superradiance cloud, are promising targets. The growth timescale which indicates how long it takes after a BH merger for EM emission to start is of order [2]
@@
corresponding to $\mathcal{O}(100)$ superradiance times, the timescale over which superradiant instabilities grow exponentially around a rotating BH. Simulations have inferred the expected electromagnetic luminosity in terms of the BH parameters and the properties of the dark photon to scale as [2]
@@
is the final spin that saturates the BH superradiance condition in Eq. (3.4). The dark photon mass that can be probed via superradiance is limited from above as $a_{*}$ approaches $a_{*,\text{fin}}$ until eventually the BH’s spin is too small, such that its angular velocity falls below the dark photon mass, and superradiance does not occur [2]. The dark photon mass that can be probed by following up after a BH merger event is limited from below by the growth timescale – if the signal only appears after decades, the planned five year run-time of DSA-2000 is too short to capture it. In our forecasts featuring BH mergers from the last few years, we take this cutoff to be $t_{\text{growth}}<8$ years, using the estimate in Eq. (3.8).
After the growth of the superradiance cloud the electromagnetic signal persists, while the cloud depletes through gravitational wave and electromagnetic emission. When more energy in released in gravitiational waves the decay of the luminosity proceeds via a power-law in time, while in the opposite scenario the decay is exponential, such that [2]
$\frac{L_{\text{EM}}(t_{\text{growth}}+t)}{L_{\text{\text{EM}}}(t_{\text{growth}})}=\begin{cases}\left(1+t/\tau_{GW}\right)^{-1}&\,\tau_{\text{GW}}\ll\tau_{EM}\,,\\ e^{(-t/\tau_{\text{EM}})\ln 2}&\tau_{GW}\gg\tau_{\text{EM}}\,,\end{cases}$ (3.11)
where the respective lifetimes scale as [2]
@@
In Table 4 we have identified several promising BH-BH merger events and NS-BH merger events with good spatial localization within the DSA-2000 field of view that could have lead to dark photon superradiance in the final BH [162, 163]. Many of these events have good spatial localization due to multi-messenger signals detected shortly after the merger event, which are particularly expected for NS-BH mergers [164, 165, 166]. The predicted signal from superradiance is distinct from these multi-messenger detections due to the predicted time delay from the initial merger that is needed for the superradiance cloud to grow and produce electromagnetic signals. We expect the O5 Observing Runs from the LIGO/VIRGO/KAGRA collaboration starting in 2028 [167] to identify several more promising targets relevant for a search with the DSA-2000 beyond the already discovered ones listed in Table 4. The spatial localization will also improve in these future runs, which is particularly promising for spatially correlating the described radio signal from superradiance with BH merger events.
Figure 8: Projected discovery potential (SNR = 5) of the continuous radio signal from a dark photon BH superradiance cloud with mixing parameter $\epsilon$ and dark photon mass $m_{\gamma^{\prime}}$ for 15 min intervals of observations as part of the all-sky survey with DSA-2000. The lower mass range is bounded by the requirement that the superradiance cloud forms on a timescale shorter than 8 years. The gray shaded region is excluded by measurements of spectral distortions in the CMB performed by COBE/FIRAS [146, 145], and Planck CMB and unWise galaxy survey data cross-correlation [147, 148, 149]. Limits on dark photon DM extend to $\epsilon\gtrsim 10^{-14}$ and are not shown [110].
@@
In Figure 8 we show the discovery reach ($\text{SNR}\equiv F_{\text{obs}}/\sigma=5$) in kinetic mixing $\epsilon$ that DSA-2000 will achieve through correlating detected broadband radio sources in its all sky survey with the spatial locations of the compact object mergers. For the shown forecast, we evaluate $L_{\text{EM}}$ at $t=t_{\text{growth}}$ in Eq. (3.9) and Eq. (3.11), and take $\eta_{\text{radio}}=10^{-4}$ in Eq. (3.14). This choice for $\eta_{\text{radio}}$ is motivated by the emission spectrum of standard pulsars [161, 71, 2]. We have also assumed that, as part of DSA-2000’s all sky survey, a sensitivity of $\sigma=2\mu$Jy is reached for every sky patch within the field of view in the 15 minute intervals of observation. The detection of positive correlations would motivate targeted follow-ups with the DSA-2000 with increased frequency resolution that can resolve the spinning up pulsing frequency of the superradiance cloud radio signal. Figure 8 shows that BH superradiance radio signals can exceed the current dark photon limits by about one order of magnitude in sensitivity up to couplings as small as $\epsilon\gtrsim 3\times 10^{-8}$ for dark photon masses between $6\times 10^{-14}\,\text{eV}-5\times 10^{-13}\,\text{eV}$. However, we note that Figure 8 shows the potential of a discovery search, rather than a projection for a robust constraint. Due to the uncertainty in the spectral composition and total fraction of the luminosity emitted in radio frequencies, quantified by $\eta_{\text{radio}}$, the absence of a signal does not rule out a dark photon in the shown mass range. However, its distinguishing features, bright radio sources spatially correlated with compact merger events turning on a timescale of months to years after the merger, in combination with a spinning up pulsing component with a pulsing frequency clustered around the dark photon mass, make it an exciting signal for a discovery search.

### v2

@@ 2.4 Axion clouds around neutron stars
Fig. 7 shows the projected reach for DSA-2000 observation of the Crab pulsar (dashed magenta), ATNF catalog (solid green), and the new population model (solid blue), requiring an SNR of $5$. We also include the projected reach for a $10$-hr observation of the best target predicted by our population model to be discovered by the DSA-2000 (dashed red), as well as an optimistic $1000$-hr observation of the same target (dotted red). Existing constraints in this mass range are shown gray. Leading astrophysical constraints on non-DM axion (darker gray) include the radio signal from polar cap axion conversion to photon exiting the NS [Noordhuis:2022ljw]. Leading terrestrial constraints on axion DM (lighter gray) includes ADMX [Asztalos_2010, ADMX:2018gho, ADMX:2018ogs, ADMX:2019uok, ADMX:2021mio, ADMX:2021nhd, ADMX:2024xbv], CAPP [Lee:2020cfj, Jeong:2020cwz, CAPP:2020utb, Lee:2022mnc, Yoon:2022gzp, Kim:2022hmg, Yi:2022fmn, Yang:2023yry, Kim:2023vpo, CAPP:2024dtx], and RBF+UF [PhysRevLett.59.839, Wuensch:1989sa, PhysRevD.42.1297, Hagmann:1996qd]. Our projections indicate that DSA-2000 will reach a sensitivity of order $g_{a\gamma\gamma}\sim 10^{-14}\,\text{GeV}^{-1}$, which will set the most competitive non-DM axion constraints in the mass range $2.9\,\mu\mathrm{eV}\lesssim m_{a}\lesssim 8.3\,\mu\mathrm{eV}$, including the uncovered parameter space in the mass range $5\,\mu\mathrm{eV}\lesssim m_{a}\lesssim 8\,\mu\mathrm{eV}$ in the axion DM landscape. This can be achieved with either a $10$-hr fixed-target search pointing at the Crab pulsar, or data from the planned all-sky survey assuming discovery of the expected new pulsars. Independent of the discovery of new pulsars, data from the planned all-sky survey will allow us to set the most competitive non-DM axion bound at $\sim 10^{-13}\,\text{GeV}^{-1}$ in the mass range $3\,\mu\mathrm{eV}\lesssim m_{a}\lesssim 8\,\mu\mathrm{eV}$, using the ATNF catalog. If DSA-2000 is able to discover a pulsar with parameters comparable to the best predicted target (see Table 3), a $10$-hr dedicated observation would allow us to reach $g_{a\gamma\gamma}\approx 5\times 10^{-15}\,\text{GeV}^{-1}$, and a $1000$-hr dedicated observation would probe the KSVZ axion hypothesis.
@@ 3 Radio signals from dark photons
Many of the observational signatures of axions are shared by another well-motivated DM candidate, the dark photon (sometimes referred to as the “hidden photon”) [Jaeckel_2010, jaeckel2013forcestandardmodel, Fabbrichesi_2021, Caputo_2021]. Indeed, resonance conversion into radio photons near NSs is also possible for dark photon DM [Hardy:2022ufh], in close analogy with the axion DM case discussed in Sec. 2. However, the limits one could obtain from such radio signals in the DSA-2000 frequency band are not competitive. Here we focus on using the DSA-2000 to probe radio signals generated through dark photon superradiance [Siemonsen:2022ivj, Mirasola:2025car], with sensitivity to dark photon masses that range from $5\times 10^{-14}-10^{-12}\,\text{eV}$.
The dark photon, $\gamma^{\prime}$, is a hypothetical vector particle with mass $m_{\gamma^{\prime}}$ that interacts with the Standard Model via kinetic mixing with ordinary photons. This is a well-motivated extension to the Standard Model that could, for example, arise from a new $U^{\prime}(1)$ gauge group. The dark photon itself is a viable DM candidate, but could also constitute a vector portal to a dark BSM sector which contain the stable DM. It is extensively searched for using laboratory probes, for example, fifth force [PhysRevLett.61.2285, Kroff:2020zhp] and light shining through wall experiments [Ehret:2010mh, Tang:2023oid], astrophysical probes [Schwarz:2015lqa, Linden:2024uph], as well as cosmological probes such as cosmic microwave background measurements (CMB)[Mirizzi:2009iz, McDermott:2019lch, Caputo:2020bdy]. Its interaction with the SM is described by the Lagrangian
@@
where $A^{\mu}$ denotes the SM photon, $F^{\mu\nu}$ is the electromagnetic field strength, $A^{\prime\mu}$ is the dark photon, $F^{\prime\mu\nu}$ is the dark photon field strength, $\epsilon$ is the kinetic mixing parameter, and $J^{\mu}$ ($J^{\prime\mu}$) is the electromagnetic (dark) current. While a priori the mixing parameter $\epsilon$ could be order one, experimental probes constrain it to be much smaller. Dark photons can oscillate during propagation in vacuum or matter from their sterile state to visible photons. This oscillation can be resonantly enhanced if the dark photon mass, $m_{\gamma^{\prime}}$, coincides with the plasma mass of the visible photon, which scales with the electron number density. This process, analogous to resonant axion-photon conversion in plasma, sets the most stringent constraints on non-DM dark photons, $\epsilon\lesssim 5\times 10^{-7}-5\times 10^{-8}$ on dark photons in the relevant mass range from $10^{-14}-10^{-11}\,\text{eV}$ from measurements of spectral distortions [Fixsen:1996nj, Caputo:2020bdy] and spectral anisotropies [Planck:2018nkj, unwise, McCarthy:2024ozh] in the CMB. [^10] We disregard the constraints from [Cardoso:2018tly] who argue that black hole superradiance rules out dark massive vector particles in the mass range from $10^{-13}-3\times 10^{-12}\,\text{eV}$ based on the observed stability of the black holes’ inner disk. Non-gravitational interactions of the cloud perturb the disk dynamics as explored in [Siemonsen:2022ivj] which invalidates an extrapolation of the result of [Cardoso:2018tly] to the finite values of $\epsilon$ we consider.
@@ 3.1 Kinetically mixed dark photon superradiance
The phenomenon of Black Hole (BH) superradiance describes the formation of a bosonic cloud around a spinning black hole [Zeldovich:1971abc, PhysRevLett.28.994, Starobinsky:1973aij, PhysRevD.22.2323, Bekenstein:1998nt, Brito:2015oca]. The bosonic cloud is a coherent gravitationally-bound state of ultralight bosons, which forms if an ultralight boson exists whose Compton wavelength is of the order of the BH horizon size [PhysRevD.81.123530, Arvanitaki:2010sy, Brito:2015oca]. This phenomenon is independent of any initial abundance such that superradiance occurs even if the boson does not comprise any sub-component of DM.
The dark photon superradiance signatures described below rely on theoretical modeling that involves significant uncertainties, particularly in the plasma physics and energy dissipation mechanisms. The theoretical framework we utilize, developed in [Siemonsen:2022ivj], makes several assumptions about: cascade pair production, resistive magnetohydrodynamic plasma modeling adapted from pulsar studies, and turbulent energy dissipation processes, which require further validation in the extreme electromagnetic environments around black holes. Given these theoretical uncertainties, we focus on discovery-oriented searches rather than setting exclusion limits, as the absence of predicted signals could result from the breakdown of theoretical assumptions rather than the non-existence of dark photons.
For a dark photon, the superradiance mechanism sources a dark photon superradiance cloud around a spinning BH [Baryakhtar:2017ngi], which if it possesses a large enough kinetic mixing with the SM photon, induces a rotating visible electromagnetic field, accompanied by a plasma of charged particles [Siemonsen:2022ivj]. The number of dark photons in the cloud grows exponentially, following the birth of a BH with initial dimensionless spin $a_{*}$, mass $M$, and angular velocity [^11] The angular velocity of a BH indicates how fast its event horizon is rotating.[Arvanitaki:2014wva, Siemonsen:2022ivj]
@@
Here $G$ denotes Newton’s gravitational constant, and [Siemonsen:2022ivj]
@@
The conducting plasma within the superradiance cloud is seeded by initial synchrotron radiation from environmental electrons accelerated in the EM field of the superradiance cloud [Siemonsen:2022ivj]. They induce photon-assisted Schwinger pair production, efficiently populating an EM plasma for kinetic mixing parameters that exceed about [Siemonsen:2022ivj]
@@
This astrophysical environment resembles a pulsar magnetosphere with distinguishing “smoking gun” features. In particular, the electromagnetic luminosity can be as high as $\sim 10^{42}\,\text{erg/s}$ [Siemonsen:2022ivj], and is expected to have a continuous flux component in the radio frequency band of $\mathcal{O}(10^{-4})$ of the total luminosity [Kaspi:2017fwg, Manchester:2004bp, Siemonsen:2022ivj], resulting in a broadband radio flux ($\mathcal{B}\sim\text{GHz}$) that could be as large as
@@
Unlike normal pulsar signals, the time derivative of the pulsing frequency is positive, indicating a spin-up, [Siemonsen:2022ivj]
@@
Since the pulsing frequency at leading order is determined by the dark photon mass, identifying several objects with these characteristics clustered around the same frequency would be a smoking gun for detecting a kinetically mixed dark photon superradiance cloud, particularly if the signals can be spatially correlated with prior BH merger events. The characteristic dark photon masses that would lead to this observable superradiance phenomenon are around $m_{\gamma^{\prime}}\sim 10^{-12}\,\text{eV}$ for stellar BHs [Siemonsen:2022ivj], which corresponds to dark photons whose Compton wavelengths are of the order of the observed BH horizon sizes.
DSA–2000 can probe this dark photon signal through different search strategies. In particular, electromagnetic follow-ups of compact binary mergers, with characteristics that allow for a timely formation ($t_{\text{growth}}\lesssim\mathcal{O}\,(\text{years})$) of the superradiance cloud, are promising targets. The growth timescale which indicates how long it takes after a BH merger for EM emission to start is of order [Siemonsen:2022ivj]
@@
corresponding to $\mathcal{O}(100)$ superradiance times, the timescale over which superradiant instabilities grow exponentially around a rotating BH. Simulations have inferred the expected electromagnetic luminosity in terms of the BH parameters and the properties of the dark photon to scale as [Siemonsen:2022ivj]
@@
is the final spin that saturates the BH superradiance condition in Eq. (3.4). The dark photon mass that can be probed via superradiance is limited from above as $a_{*}$ approaches $a_{*,\text{fin}}$ until eventually the BH’s spin is too small, such that its angular velocity falls below the dark photon mass, and superradiance does not occur [Siemonsen:2022ivj]. The dark photon mass that can be probed by following up after a BH merger event is limited from below by the growth timescale – if the signal only appears after decades, the planned five year run-time of DSA-2000 is too short to capture it. In our forecasts featuring BH mergers from the last few years, we take this cutoff to be $t_{\text{growth}}<8$ years, using the estimate in Eq. (3.8).
After the growth of the superradiance cloud the electromagnetic signal persists, while the cloud depletes through gravitational wave and electromagnetic emission. When more energy in released in gravitiational waves the decay of the luminosity proceeds via a power-law in time, while in the opposite scenario the decay is exponential, such that [Siemonsen:2022ivj]
[…]
where the respective lifetimes scale as [Siemonsen:2022ivj]
@@
In Table 4 we have identified several promising BH-BH merger events and NS-BH merger events with good spatial localization within the DSA-2000 field of view that could have lead to dark photon superradiance in the final BH [LIGOScientific:2018mvr, KAGRA:2021vkt]. Many of these events have good spatial localization due to multi-messenger signals detected shortly after the merger event, which are particularly expected for NS-BH mergers [2019ApJ...887L..13D, 2021ApJ...923...66A, Dobie:2021khu]. The predicted signal from superradiance is distinct from these multi-messenger detections due to the predicted time delay from the initial merger that is needed for the superradiance cloud to grow and produce electromagnetic signals. We expect the O5 Observing Runs from the LIGO/VIRGO/KAGRA collaboration starting in 2028 [KAGRA:2013rdx] to identify several more promising targets relevant for a search with the DSA-2000 beyond the already discovered ones listed in Table 4. The spatial localization will also improve in these future runs, which is particularly promising for spatially correlating the described radio signal from superradiance with BH merger events.
Figure 8: Projected discovery potential (SNR = 5) of the continuous radio signal from a dark photon BH superradiance cloud with mixing parameter $\epsilon$ and dark photon mass $m_{\gamma^{\prime}}$ for 15 min intervals of observations as part of the all-sky survey with DSA-2000. The lower mass range is bounded by the requirement that the superradiance cloud forms on a timescale shorter than 8 years. The gray shaded region is excluded by measurements of spectral distortions in the CMB performed by COBE/FIRAS [Fixsen:1996nj, Caputo:2020bdy], and Planck CMB and unWise galaxy survey data cross-correlation [Planck:2018nkj, unwise, McCarthy:2024ozh]. Limits on dark photon DM extend to $\epsilon\gtrsim 10^{-14}$ and are not shown [AxionLimits].
@@
[…]